OpenStudy (anonymous):

Can you help me Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). A: y − 4 = −one half(x + 6) B: y − 4 = 2(x + 6) C: y + 6 = −one half(x − 4) D: y + 6 = 2(x − 4)

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

first we need to find the slope by using the slope formula: slope(m) = (y2 - y1) / (x2 - x1) (6,4) x1 = 6, y1 = 4 (2,0) x2 = 2, y2 = 0 now lets sub our info into the formula m = (0 - 4) / (2 - 6) m = -4/-4 m = 1 understand so far ?

OpenStudy (anonymous):

i did the math i just wanna make sure that is correct because im still getting use to working these problems out

OpenStudy (anonymous):

I messed up on the slope...it is -4/8 or -1/2 y - y1 = m(x - x1) using slope -1/2 and points (-6,4) y - 4 = -1/2(x - (-6) y - 4 = -1/2(x + 6) your answer is correct

OpenStudy (anonymous):

thank u soooo @kelliegirl33

OpenStudy (anonymous):

no problem :)